The polynomial f(x) divided x – 3 results in a quotient of x2 + 3x – 5
with a remainder of 2. Find f(3).

Answer by pvrbrahmam(12)

Q: The polynomial f(x) divided x – 3 results in a quotient of x2 + 3x – 5
with a remainder of 2. Find f(3).
Sol : f(x) = quotient *Divisor + remainder
= (x^2 +3x-5)(x-3) + 2
f(3) = (3^2 + 3*3-5)(3-3)+2= 2